gokussgss20:

Sylvie finds the solution to the system of equations by graphing. y = A system of equations. y equals StartFraction 2 over 3 EndFraction x plus 1. y equals negative StartFraction 2 over 3 EndFraction x minus 1x + 1 and y = x – 1 Which graph shows the solution to Sylvie’s system of equations?

Mercury:

y = (2/3)x + 1 y = -(2/3)x -x + 1? (might be wrong on this?) y = x - 1 anyway, you just need to find the graph that correctly shows all three equations, and marks their intersection as the solution. check the slope and y -intercept of each graph to make sure it matches.